# Ferraz, Cappart, and Vidal

### v1

@@ 6.1 Data Collection and Experimental Setup
Geographical data and test instances. We base our studies on five metropolitan areas located in the UK (London, Bristol, Manchester, Leeds, and West-Midlands). The selection of these areas was driven by their diversity and availability of precise geographical boundaries from https://movement.uber.com as well as population statistics from the UK government 2018 census database (Park 2018). The BUs correspond to Middle Layer Super Output Areas (MSOAs), which are designed to contain roughly the same population (each MSOA contains over 5,000 inhabitants, and 8,000 inhabitants on average over the UK). Finally, to obtain data sets with a different number of BUs ($n=\{60,90,120\}$), we selected a center point in each region and retained the $n$ closest BUs.
Table 1 provides general statistics (population, area, density, and compactness) on the $n=120$ BUs of each metropolitan area. The area of each BU was measured using the Monte Carlo method with 50,000 samples. Compactness scores have been obtained using Reocks formula (Young 1988), which divides the area of the BU by the area of the smallest circumscribed circle. Therefore, this compactness measure assigns higher scores to areas that are more circularly shaped.
@@
Population (thousands) average
8.32
7.53
9.68
8.69 8.71
std
2.29
1.64
2.03
2.36 2.06
min
5.55
5.20
6.58
5.26 5.44
median
7.68
7.26
9.40
8.32 8.19
max
18.16
14.06
16.17
15.87 17.12
[…]
10.34
4.67
0.75
2.22 1.81
std
22.42
7.22
0.47
1.17 0.75
min
0.63
0.35
0.30
0.59 0.53
median
1.99
2.51
0.64
1.90 1.68
max
171.21
51.79
3.58
6.69 4.39
[…]
3.92
3.61
15.15
4.87 5.51
std
2.89
3.41
4.91
2.58 2.41
min
0.06
0.14
2.76
1.12 1.99
median
3.74
2.90
15.17
4.41 5.21
max
12.72
25.20
28.27
16.36 17.06
Compactness average
0.40
0.43
0.43
0.42 0.44
std
0.10
0.09
0.09
0.10 0.09
min
0.19
0.25
0.21
0.20 0.22
median
0.40
0.43
0.43
0.42 0.44
max
0.65
0.67
0.63
0.66 0.61
Table 1: Population and geography statistics for the BUs
[…]
Besides the characteristics of the BUs, other factors impact the structure of the district-and-routing problems: the admissible range for the number of BUs in each district [$n_{\textsc{l}},n_{\textsc{u}}]$, and the location of the depot. We generated five different configurations for the district-size constraints by setting $n_{\textsc{l}}=\lfloor 0.8\times t\rfloor$ and $n_{\textsc{u}}=\lceil 1.2\times t\rceil$ with $t\in\{3,6,12,20,30\}$. Then, for each configuration, the probability $\kappa$ that an inhabitant will make a request has been set to $\kappa=96/(8000t)$ where $96$ corresponds to the targeted number of requests in the related routing problem, and $8000$ corresponds to the average population of a BU. The values for parameter $\kappa$ have been selected to reflect realistic scenarios with delivery routes that cover close to a hundred stops (typical for parcel deliveries). Finally, five possible locations for the depot were considered: at the center of the metropolitan area (C), at the north-east (NE), at the north-west (NW), at the south-east (SE), and at the south-west (SW) of the center ($D=\{\texttt{C},\texttt{NE},\texttt{NW},\texttt{SE},\texttt{SW}\}$). All these factors are summarized in Table 2. For each of the five considered metropolitan areas, we therefore generated $3\times 5\times 5=75$ instances covering all the possible combinations of these factors.
@@
[…]
Minimum number of BUs in a district $\lfloor 0.8\times t\rfloor$
Maximum number of BUs in a district $\lceil 1.2\times t\rceil$
@@
Table 2: Summary of the different instance parameters
Scenarios and solution evaluations. The districting-and-routing problem has an objective function that must be measured over a large number of scenarios. Consequently, to standardize the evaluation procedures among the different evaluation and solution methods, it was essential to (i) create separate demand scenarios (positions of customers’ demands) for training and testing and (ii) adopt a standard for solution measurement. We did this by sampling for each BU $n_{\textsc{Train}}=500$ and $n_{\textsc{Test}}=500$ random scenarios for training and testing, respectively. Therefore, let $\mathcal{R}^{\textsc{x}}_{it}$ be the set of customer requests, characterized by their positions, for any BU $i$ and scenario $t$ for $\textsc{x}\in\{\textsc{Train},\textsc{Test}\}$. All methods have only access to the training set of scenarios during the learning and solution process. We reserve the other scenarios for the final evaluation and comparison of the solutions. For each demand scenario of the BUs, we obtain a corresponding demand scenario for any district by compounding all the demand locations of the BUs it contains. With this, the training and testing cost of a solution $\mathcal{S}$ described as a set of districts are defined as the average cost of the TSP tours over the scenarios, calculated as:
$\Phi^{\textsc{x}}_{\textsc{saa}}(\mathcal{S})=\frac{1}{n_{\textsc{x}}}\sum_{t=1}^{n_{\textsc{x}}}\sum_{d\in\mathcal{S}}C_{\textsc{tsp}}\left(\bigcup_{i\in d}\mathcal{R}^{\textsc{x}}_{it}\right),$ (19)
for $\textsc{x}\in\{\textsc{Train},\textsc{Test}\}$, and where $C_{\textsc{tsp}}(\mathcal{R})$ calculates the cost of a TSP tour visiting the depot and all customers in $\mathcal{R}$. We rely on Lin-Kernighan (LKH) algorithm (from Helsgaun 2000, available at http://webhotel4.ruc.dk/~keld/research/LKH/) to measure TSP distances. The reason for this choice is that this algorithm is very fast (0.1 seconds per run) and finds almost systematically the optimal solution in our cases with roughly 100 customer requests, so there is no need to employ time-consuming exact approaches that would limit our experimentation capabilities.
Computational Environment. The calculation of TSP scenarios to estimate district costs has been conducted on Intel Gold 6148 Skylake 2.4 GHz with 80GB RAM. The rest of the experimental analyses (including training, calibration, and optimization) have been conducted on a computer with an Intel E5-2683 v4 Broadwell 2.1GHz CPU, 124GB RAM, and an Nvidia P100 Pascal (12GB HBM2 memory) GPU. To implement the GNN, we use the official implementation of structure2vec through its Python interface, available at https://github.com/Hanjun-Dai/pytorch_structure2vec. The ILS is implemented in C++, compiled with G++ v9.3.0. Finally, for experimental reproducibility, all the data, scripts, and source code needed to reproduce our experiments are provided in open source at https://github.com/vidalt/Districting-Routing.
@@ 6.2 Estimation Methods for Routing Costs
  • BHHD: Let $d$ be a district composed of $n$ BUs. For any BU $i$, we recall that $a_{i}$ is the area of the BU and $\xi_{i}$ is its population. Moreover, $\kappa$ is the probability of a delivery request for any inhabitant. With this, the total area of the district is calculated as $A_{d}=\sum_{i\in d}a_{i}$, and the expected number of deliveries within the district is $R_{d}=\kappa\sum_{i\in d}\xi_{i}$. Extending the BHH formula to account for the distance from the depot as in Daganzo (1984) gives:
    $\Phi_{\textsc{bd}}(d)=\beta\sqrt{A_{d}R_{d}}+2\Delta_{d},$ (20)
    where $\beta$ is a hyper-parameter that needs calibration, and $\Delta_{d}$ is the average distance between the depot and a request, calculated by Monte-Carlo estimation on the training scenario set. In our experiments, we set $\beta$ to minimize the mean-squared error over a training set (discussed at the end of this section) through least-squares regression. We note that BHH formula converges almost surely to the true expected distance in an asymptotic regime where the number of delivery requests tends towards infinity. As a consequence, approximation errors naturally occur as we deviate from this assumption.
  • FIG: This continuous approximation formula of Figliozzi (2007) is a direct extension of Equation (20), which was designed to cope with the practical non-uniformity of real observed demands. The formula is defined in Equation (21), and includes four hyper-parameters ($\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$) requiring calibration. As previously, $\Delta_{d}$ represents the average distance between the depot and the customer requests. As previously, the values of the hyperparameters are selected through least-squares regression.
    $\texttt{FIG}(d)=\beta_{1}\sqrt{A_{d}R_{d}}+\beta_{2}\Delta_{d}+\beta_{3}\sqrt{\frac{A_{d}}{R_{d}}}+\beta_{4}\$ (21)
  • SNN: This estimator is based on a neural network with one hidden layer of three neurons. Five features are used as input of the network:
@@
    Kwon, Golden, and Wasil (1995) relied on the sigmoid function for activation and used standard backpropagation for training. However, it is noteworthy that this architecture was designed in 1995 and consequently did not leverage the extensive recent improvements of deep learning. Consequently, our preliminary experiments showed that this network performed poorly compared to the other baselines. To obtain a fair comparison leveraging newer training strategies, we decided to rely on ReLU activation functions (instead of sigmoid) and carry out training using Adam optimizer (Kingma and Ba 2014) with a learning rate of $10^{-3}$. The SNN model is therefore trained for $50,000$ epochs considering the mean square error as loss function.
Calibration and Training. To calibrate and train the different cost-estimation methods, we sampled for each instance 9,000 random connected districts respecting the size constraints $[n_{\textsc{l}},n_{\textsc{u}}]$. For each of these districts, we calculated the expected TSP cost by SAA on the Train demand scenarios. For GNN and SNN, we further subdivided this set into 8,000 districts for training and 1,000 districts for validation to control convergence. Training the GNN takes two hours on average, i.e., a substantial but reasonable computational effort since strategic districting decisions hold for a long time.
6.3 Results – Prediction Accuracy
Our first set of experiments aims to evaluate the accuracy of the different models (BHHD, FIG, SNN, as well as the proposed GNN) for estimating the routing costs. Therefore, we use the trained models described in the previous section and evaluate them on an additional set of 1,000 random districts that are distinct from the ones used during training. We compare the estimates with “ground truth” values of the routing costs obtained again by SAA over the $500$ Test demand scenarios. In the rest of this section, we analyze how the proposed estimation approaches deviate from the ground-truth measurements in terms of their root-mean-square error (RMSE – the lower, the better). For a given test instance, the RMSE is calculated as:
$\text{RMSE}=\sqrt{\frac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\left(\Phi(d)-\Phi^{\textsc{test}}_{\textsc{saa}}(d)\right)^{2}},$ (23)
@@ 6.3 Results – Prediction Accuracy
$\bm{n}$ $\bm{t}$ $\hat{\Phi}^{\textsc{test}}_{\textsc{saa}}$ RMSE of Estimation Method
GNN BHHD FIG SNN
60 3 43.16 $\pm$ 0.48 1.48 2.27 2.09 2.05
6 48.21 $\pm$ 0.66 1.39 3.43 3.24 3.02
12 52.06 $\pm$ 0.80 2.10 4.36 4.15 4.00
20 56.42 $\pm$ 0.95 2.39 5.18 4.89 4.73
30 61.14 $\pm$ 1.04 2.08 4.73 4.50 4.31
90 3 52.43 $\pm$ 0.53 1.56 2.70 2.49 2.47
6 57.03 $\pm$ 0.69 1.88 4.00 3.75 3.60
12 61.19 $\pm$ 0.83 2.62 5.12 4.80 4.60
20 66.67 $\pm$ 1.02 3.00 6.17 5.74 5.50
30 72.63 $\pm$ 1.14 2.84 6.15 5.78 5.50
120 3 61.61 $\pm$ 0.60 1.68 3.27 3.15 3.10
6 63.25 $\pm$ 0.69 2.09 4.04 3.88 3.76
12 67.57 $\pm$ 0.84 2.83 5.28 5.00 4.84
20 74.44 $\pm$ 1.07 3.34 6.36 6.03 5.90
30 80.81 $\pm$ 1.21 3.39 6.82 6.52 6.21
Average 59.84 $\pm$ 0.81 2.31 4.66 4.40 4.24
@@
Depot Metropolitan Area $\hat{\Phi}^{\textsc{test}}_{\textsc{saa}}$ RMSE of Estimation Method
GNN BHHD FIG SNN
[…]
Leeds 48.71 $\pm$ 0.97 3.14 5.75 5.25 5.29
London 28.58 $\pm$ 0.59 1.75 3.25 3.19 3.11
Manchester 43.13 $\pm$ 0.89 2.70 4.51 4.43 4.45
West-Midlands 40.59 $\pm$ 0.84 2.24 4.06 3.90 3.88
[…]
Leeds 71.78 $\pm$ 0.86 2.67 5.70 5.09 5.26
London 37.03 $\pm$ 0.54 1.69 2.98 2.93 2.90
Manchester 57.60 $\pm$ 0.80 2.56 4.14 4.09 4.11
West-Midlands 53.39 $\pm$0.76 2.30 3.70 3.66 3.62
@@
As seen in these tables, with an overall RMSE of 2.31, the quality of the estimate obtained with the proposed GNN is vastly superior to that of all the other methods: BHHD with a RMSE of 4.66, FIG with a RMSE of 4.40, and SNN with a RMSE of 4.24. The error committed also appears to increase with the number of BUs in the metropolitan area ($n$) as well as the target number of BUs ($t$) in each district. This is due to two factors. Firstly, the magnitude of the predicted values (long-term operational cost of the districts) grows with $t$, and also with $n$ to a lesser extent (instances with larger $n$ include less-populated BUs that are located farther away from city centers). Therefore, as these values grow larger, more estimation error is generally committed. Secondly, the number of possible districts grows exponentially in $n$ and $t$, such that the universe of possible inputs quickly grows, and it becomes harder to learn the target.
@@
($n=90$, $t=3$)
($n=90$, $t=12$)
($n=120$, $t=12$)

### v2

@@ 6.1 Data Collection and Experimental Setup
Geographical data and test instances. We base our studies on five metropolitan areas located in the UK (London, Bristol, Manchester, Leeds, and the West Midlands). The selection of these areas was driven by their diversity and availability of precise geographical boundaries from https://movement.uber.com as well as population statistics from the UK government 2018 census database (Park 2018). The BUs correspond to Middle Layer Super Output Areas (MSOAs), which are designed to contain roughly the same population (each MSOA contains over 5,000 inhabitants, and 8,000 inhabitants on average over the UK). Finally, to obtain data sets with a different number of BUs ($n=\{60,90,120\}$), we selected a center point in each region and retained the $n$ closest BUs.
Table 1 provides general statistics (population, area, density, and compactness) on the $n=120$ BUs of each metropolitan area. The area of each BU was measured using the Monte Carlo method with 50,000 samples. Compactness scores have been obtained using Reock’s formula (Young 1988), which divides the area of the BU by the area of the smallest circumscribed circle. Therefore, this compactness measure assigns higher scores to areas that are more circularly shaped.
Table 1: Population and geography statistics for the BUs
@@
Population (thousands) average 8.32 7.53 9.68 8.69 8.71
std 2.29 1.64 2.03 2.36 2.06
min 5.55 5.20 6.58 5.26 5.44
median 7.68 7.26 9.40 8.32 8.19
max 18.16 14.06 16.17 15.87 17.12
Area (km2) average 10.34 4.67 0.75 2.22 1.81
std 22.42 7.22 0.47 1.17 0.75
min 0.63 0.35 0.30 0.59 0.53
median 1.99 2.51 0.64 1.90 1.68
max 171.21 51.79 3.58 6.69 4.39
Density (thousands/km2) average 3.92 3.61 15.15 4.87 5.51
std 2.89 3.41 4.91 2.58 2.41
min 0.06 0.14 2.76 1.12 1.99
median 3.74 2.90 15.17 4.41 5.21
max 12.72 25.20 28.27 16.36 17.06
Compactness average 0.40 0.43 0.43 0.42 0.44
std 0.10 0.09 0.09 0.10 0.09
min 0.19 0.25 0.21 0.20 0.22
median 0.40 0.43 0.43 0.42 0.44
max 0.65 0.67 0.63 0.66 0.61
As seen in Table 1, there are important differences in geography and population between the BUs of the five metropolitan areas. Bristol and Leeds BUs generally have smaller population densities and larger geographical extensions, whereas London has the densest and smallest BUs. BUs areas are also very disparate, ranging from 0.30 km2 (in London) to 171.21 km2 (in Bristol). Finally, compactness measures appear to be roughly similar for the five cities.
Besides the characteristics of the BUs, other factors impact the structure of the district-and-routing problems: the admissible range for the number of BUs in each district $n_{\textsc{l}},n_{\textsc{u}}]$, and the location of the depot. For $t\in\{3,6,12,20,30\}$, we generated five different configurations for the district-size constraints by setting $n_{\textsc{l}}=\lfloor 0.8\times t\rfloor$ and $n_{\textsc{u}}=\lceil 1.2\times t\rceil$, and setting the number of districts to $k=\lfloor n/t\rfloor$. Then, for each configuration, the probability $\kappa$ that an inhabitant will make a request has been set to $\kappa=96/(8000t)$ where $96$ corresponds to the targeted number of requests in the related routing problem, and $8000$ corresponds to the average population of a BU. The values for parameter $\kappa$ have been selected to reflect realistic scenarios with delivery routes that cover close to a hundred stops (typical for parcel deliveries, as seen in Holland et al. 2017). Finally, five possible locations for the depot were considered: at the center of the metropolitan area (C), at the north-east (NE), at the north-west (NW), at the south-east (SE), and at the south-west (SW) of the center ($D=\{\texttt{C},\texttt{NE},\texttt{NW},\texttt{SE},\texttt{SW}\}$). All these factors are summarized in Table 2. For each of the five considered metropolitan areas, we therefore generated $3\times 5\times 5=75$ instances covering all the possible combinations of these factors.
Table 2: Summary of the different instance parameters
@@
Depot location $\{\texttt{C},\texttt{NE},\texttt{NW},\texttt{SE},\texttt{SW}\}$
Minimum number of BUs in a district $n_{\textsc{l}}=\lfloor 0.8\times t\rfloor$
Maximum number of BUs in a district $n_{\textsc{u}}=\lceil 1.2\times t\rceil$
Number of districts $k=\lfloor n/t\rfloor$
@@
Scenarios and solution evaluations. The districting-and-routing problem has an objective function that must be measured over a large number of scenarios. Consequently, to standardize the evaluation procedures among the different evaluation and solution methods, it was essential to (i) create separate demand scenarios (positions of customers’ demands) for training and testing and (ii) adopt a standard for solution measurement. We did this by sampling for each BU $S_{\textsc{Train}}=500$ and $S_{\textsc{Test}}=500$ random scenarios for training and testing, respectively. Let $X^{\textsc{T}}_{it}$ be the set of customer requests, characterized by their positions, for any BU $i$ and scenario $s$ for $\textsc{T}\in\{\textsc{Train},\textsc{Test}\}$. All methods have access only to the training set of scenarios during the learning and solution process. We reserve the other scenarios for the final evaluation and comparison of the solutions. For each demand scenario of the BUs, we obtain a corresponding demand scenario for any district by compounding all the demand locations of the BUs it contains. With this, the training (and respectively testing) cost of a solution $\mathcal{S}$ described as a set of districts, is defined as the average cost of the TSP tours over the scenarios, calculated as:
$\Phi^{\textsc{T}}_{\textsc{saa}}(\mathcal{S})=\frac{1}{S_{\textsc{T}}}\sum_{s=1}^{S_{\textsc{t}}}\sum_{d\in\mathcal{S}}C_{\textsc{tsp}}\left(\bigcup_{i\in d}X^{\textsc{T}}_{is}\right),$ (19)
for $\textsc{T}\in\{\textsc{Train},\textsc{Test}\}$, and where $C_{\textsc{tsp}}(X)$ calculates the cost of a TSP tour visiting the depot and all customers in $X$. We rely on the Lin-Kernighan (LKH) algorithm (from Helsgaun 2000, available at http://webhotel4.ruc.dk/~keld/research/LKH/) to measure TSP distances. The reason for this choice is that this algorithm is very fast (0.1 seconds per run) and finds the optimal solution almost systematically in our cases with roughly 100 customer requests, so there is no need to employ time-consuming exact approaches that would limit our experimentation capabilities.
Computational Environment. The calculation of TSP scenarios to estimate district costs has been conducted on Intel Gold 6148 Skylake 2.4 GHz CPUs with 80GB RAM. The rest of the experimental analyses (including training, calibration, and optimization) have been conducted on a computer with an Intel E5-2683 v4 Broadwell 2.1GHz CPU, 124GB RAM, and an Nvidia P100 Pascal (12GB HBM2 memory) GPU. To implement the GNN, we use the official structure2vec implementation via its Python interface, available at https://github.com/Hanjun-Dai/pytorch_structure2vec. The ILS is implemented in C++, compiled with G++ v9.3.0. Finally, to ensure experimental reproducibility, all data, scripts, and source code needed to reproduce our experiments are available in an open-source repository at https://github.com/vidalt/Districting-Routing.
@@ 6.2 Estimation Methods for Routing Costs
  • BHHD: Let $d$ be a district composed of $n$ BUs. For any BU $i$, we recall that $a_{i}$ is the area of the BU and $\xi_{i}$ is its population. Moreover, $\kappa$ is the probability of a delivery request for any inhabitant. With this, the total area of the district is calculated as $A_{d}=\sum_{i\in d}a_{i}$, and the expected number of deliveries within the district is $R_{d}=\kappa\sum_{i\in d}\xi_{i}$. Extending the BHH formula to account for the distance from the depot, as in Daganzo (1984) gives:
    $\Phi_{\textsc{bd}}(d)=\alpha\sqrt{A_{d}R_{d}}+2\Delta_{d},$ (20)
    where $\alpha$ is a hyper-parameter that needs calibration, and $\Delta_{d}$ is the average distance between the depot and a request, calculated by Monte-Carlo estimation on the training scenario set. In our experiments, we set $\alpha$ to minimize the mean-squared error over a training set (discussed at the end of this section) through least-squares regression. We note that the BHH formula converges almost surely to the true expected distance in an asymptotic regime where the number of delivery requests tends towards infinity. As a consequence, approximation errors naturally occur as we deviate from this assumption.
  • FIG: This continuous approximation formula of Figliozzi (2007) is a direct extension of Equation (20), which was designed to cope with the practical non-uniformity of real observed demands. The formula is defined in Equation (21), and includes four hyper-parameters ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, $\alpha_{4}$) requiring calibration. As previously, $\Delta_{d}$ represents the average distance between the depot and the customer requests. As previously, the values of the hyperparameters are selected through least-squares regression.
    $\texttt{FIG}(d)=\alpha_{1}\sqrt{A_{d}R_{d}}+\alpha_{2}\Delta_{d}+\alpha_{3}\sqrt{\frac{A_{d}}{R_{d}}}+\alpha_{4}\\$ (21)
  • SNN: This estimator is based on a neural network with one hidden layer of three neurons. Five features are used as input to the network:
@@
    Kwon, Golden, and Wasil (1995) used a sigmoid activation function and standard backpropagation for training. However, this early architecture did not benefit from the subsequent advances in deep learning and performed poorly compared to the other baselines. To ensure a fair comparison under modern training conditions, we replaced the sigmoid activations with ReLU functions and trained the network using the Adam optimizer (Kingma and Ba 2014) with a learning rate of $10^{-3}$. The SNN model is therefore trained for $50,000$ epochs, considering the mean squared error as the loss function.
Calibration and Training. To calibrate and train the different cost-estimation methods, we sampled for each instance 9,000 random connected districts respecting the size constraints $n_{\textsc{l}},n_{\textsc{u}}]$. For each of these districts, we calculated the expected TSP cost by SAA on the Train demand scenarios. For GNN and SNN, we further subdivided this set into 8,000 districts for training and 1,000 districts for validation to control convergence. Training the GNN takes two hours on average, i.e., a substantial but reasonable computational effort since strategic districting decisions hold for a long time. Afterwards, district cost inference takes on average 8 milliseconds on a single CPU thread.
6.3 Analysis – Prediction Accuracy
Our first set of experiments aims to evaluate the accuracy of the different models (BHHD, FIG, SNN, as well as the proposed GNN) for estimating the routing costs. Therefore, we use the trained models described in the previous section and evaluate them on an additional set of 1,000 random districts that are distinct from the ones used during training. We compare the estimates with “ground truth” values of the routing costs obtained again by SAA over the $500$ Test demand scenarios. In the rest of this section, we analyze how the proposed estimation approaches deviate from the ground-truth measurements in terms of their Root-Mean-Square Error (RMSE) and Mean Absolute Percentage Error (MAPE). For a given test instance, these values are calculated as:
$\text{RMSE}=\sqrt{\frac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\left(\Phi(d)-\Phi^{\textsc{test}}_{\textsc{saa}}(d)\right)^{2}},\hskip 28.45274pt\text{MAPE}=\frac{100}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\left|\frac{\Phi(d)-\Phi^{\text{TEST}}_{\text{SAA}}(d)}{\Phi^{\text{TEST}}_{\text{SAA}}(d)}\right|,$ (23)
@@ 6.3 Analysis – Prediction Accuracy
$\bm{n}$ $\bm{t}$ $\hat{\Phi}^{\textsc{test}}_{\textsc{saa}}$ RMSE MAPE
GNN BHHD FIG SNN GNN BHHD FIG SNN
60 3 43.16 $\pm$ 0.48 1.48 2.27 2.09 2.05 2.75 3.81 3.60 3.60
6 48.21 $\pm$ 0.66 1.39 3.43 3.24 3.02 2.20 5.13 4.87 4.91
12 52.06 $\pm$ 0.80 2.10 4.36 4.15 4.00 2.85 6.19 5.78 5.79
20 56.42 $\pm$ 0.95 2.39 5.18 4.89 4.73 2.83 6.62 6.11 6.18
30 61.14 $\pm$ 1.04 2.08 4.73 4.50 4.31 2.41 5.92 5.49 5.56
90 3 52.43 $\pm$ 0.53 1.56 2.70 2.49 2.47 2.47 3.68 3.53 3.55
6 57.03 $\pm$ 0.69 1.88 4.00 3.75 3.60 2.52 4.93 4.66 4.73
12 61.19 $\pm$ 0.83 2.62 5.12 4.80 4.60 3.05 6.11 5.60 5.76
20 66.67 $\pm$ 1.02 3.00 6.17 5.74 5.50 3.00 6.75 6.12 6.27
30 72.63 $\pm$ 1.14 2.84 6.15 5.78 5.50 2.75 6.33 5.91 6.00
120 3 61.61 $\pm$ 0.60 1.68 3.27 3.15 3.10 2.27 3.60 3.46 3.49
6 63.25 $\pm$ 0.69 2.09 4.04 3.88 3.76 2.55 4.57 4.35 4.42
12 67.57 $\pm$ 0.84 2.83 5.28 5.00 4.84 3.06 5.71 5.29 5.38
20 74.44 $\pm$ 1.07 3.34 6.36 6.03 5.90 3.15 6.23 5.77 5.84
30 80.81 $\pm$ 1.21 3.39 6.82 6.52 6.21 2.91 6.25 5.78 5.88
Average 59.84 $\pm$ 0.81 2.31 4.66 4.40 4.24 2.72 5.46 5.09 5.16
@@
Depot Metropolitan Area $\hat{\Phi}^{\textsc{test}}_{\textsc{saa}}$ RMSE MAPE
GNN BHHD FIG SNN GNN BHHD FIG SNN
Central Depot $\{\texttt{C}$} Bristol 67.14 $\pm$ 1.19 4.04 8.38 7.99 6.56 4.38 8.74 9.46 7.97
Leeds 48.71 $\pm$ 0.97 3.14 5.75 5.25 5.29 4.32 8.14 8.12 7.35
London 28.58 $\pm$ 0.59 1.75 3.25 3.19 3.11 4.02 8.43 7.90 8.25
Manchester 43.13 $\pm$ 0.89 2.70 4.51 4.43 4.45 4.45 8.05 7.35 7.43
West-Midlands 40.59 $\pm$ 0.84 2.24 4.06 3.90 3.88 3.73 8.17 7.21 7.06
Off-Centered Depot $\{\texttt{NE},\texttt{NW},\texttt{SE},\texttt{SW}\}$ Bristol 105.59 $\pm$ 1.13 3.72 8.96 8.06 7.01 2.25 5.93 5.30 4.50
Leeds 71.78 $\pm$ 0.86 2.67 5.70 5.09 5.26 2.47 5.28 5.13 4.97
London 37.03 $\pm$ 0.54 1.69 2.98 2.93 2.90 3.12 5.96 5.61 5.78
Manchester 57.60 $\pm$ 0.80 2.56 4.14 4.09 4.11 3.05 5.41 5.06 5.19
West-Midlands 53.39 $\pm$0.76 2.30 3.70 3.66 3.62 2.96 5.19 4.91 5.04
As seen in these tables, with an overall RMSE of 2.31 (respectively, a MAPE of 2.72), the quality of the estimates obtained with the proposed GNN is vastly superior to that of all the other methods: BHHD with a RMSE of 4.66 (MAPE 5.46), FIG with a RMSE of 4.40 (MAPE 5.09), and SNN with a RMSE of 4.24 (MAPE 5.16). The error committed also appears to increase with the number of BUs in the metropolitan area ($n$) as well as the target number of BUs ($t$) in each district. This is due to two factors. Firstly, the magnitude of the predicted values (long-term operational cost of the districts) grows with $t$, and also with $n$ to a lesser extent (instances with larger $n$ include less-populated BUs that are located farther away from city centers). Therefore, as these values grow larger, more estimation error is generally committed. Secondly, the number of possible districts grows exponentially in $n$ and $t$, such that the universe of possible inputs quickly grows, and it becomes harder to learn the target.
@@
Finally, Figure 3 provides graphical representations of the ground-truth costs computed by SAA, along with their confidence intervals and the predicted cost values from the estimation methods. The light yellow area represents the cost provided by SAA wrapped by its 95% confidence interval. The darker continuous curve corresponds to SNN, and the lighter continuous curve corresponds to the GNN. To facilitate the visualization, we display the results for a random subset of 300 districts ordered by increasing ground-truth cost. The figure displays these graphs for two different metropolitan areas (Bristol and London), considering centered depots and three possible configurations for the $n$ and $t$ factors. Generally, we can observe that GNN provides better estimates than SNN for each scenario. Beyond this, comparing the figures with $t=3$ and $t=12$, we notice that lower $t$ values lead to much smaller errors, i.e., GNN and SNN estimates are much closer to the true cost. Larger values of $n$ also tend to diminish prediction accuracy, but this effect is less marked than when changing the district sizes.
@@
($n=90$, $t=3$) Refer to caption Refer to caption
($n=90$, $t=12$) Refer to caption Refer to caption
($n=120$, $t=12$) Refer to caption Refer to caption